I am having trouble with this question can some one please help me

I opened up a savings account today at Carthage Federal Savings and Loan with $800.00. If the yearly interest rate of 4% is compounded each month, how much interest will I earn in six months?
Thank you for your help in advance.



Answer by stanbon(75887)

I opened up a savings account today at Carthage Federal Savings and Loan with $800.00. If the yearly interest rate of 4% is compounded each month, how much interest will I earn in six months?
------------------
A(t) = P(1+(r/n))^(nt)
A(1/2) = 800(1+(0.04/12)^(12*(1/2))
A(1/2) = 800(1+0.0033)^6
A(1/2) = 800*1.02017
A(1/2) = $816.13
---------------------
Interest is $16.13
========================
Cheers,
Stan H.